# Adaptive Alpha Weighting with PPO: Enhancing Prompt-Based LLM-Generated Alphas in Quant Trading

### v1

@@ Abstract
This paper proposes a reinforcement learning framework that employs Proximal Policy Optimization (PPO) to dynamically optimize the weights of multiple large language model (LLM)-generated formulaic alphas for stock trading strategies. Formulaic alphas are mathematically defined trading signals derived from price, volume, sentiment, and other data. Although recent studies have shown that LLMs can generate diverse and effective alphas, a critical challenge lies in how to adaptively integrate them under varying market conditions. To address this gap, we leverage the deepseek-r1-distill-llama-70b model to generate fifty alphas for five major stocks: Apple, HSBC, Pepsi, Toyota, and Tencent, and then use PPO to adjust their weights in real time. Experimental results demonstrate that the PPO-optimized strategy achieves strong returns and high Sharpe ratios across most stocks, outperforming both an equal-weighted alpha portfolio and traditional benchmarks such as the Nikkei 225, S&P 500, and Hang Seng Index. The findings highlight the importance of reinforcement learning in the allocation of alpha weights and show the potential of combining LLM-generated signals with adaptive optimization for robust financial forecasting and trading.
@@ 1 Introduction
Over the past few decades, financial markets have evolved significantly with the integration of advanced technologies, especially in quantitative finance. Traditionally, stock trading strategies have relied on financial indicators such as moving averages, volatility measures, and momentum indicators to predict market behavior and generate profits 9580172 ; MOSTAFAVI2025100631 . These methods have been successful to some extent, but often lack the flexibility and adaptability required to respond to changing market conditions. In addition, relying solely on manually chosen indicators may restrict the discovery of new patterns within the extensive and constantly changing financial data landscape.
@@
Furthermore, trading signals often lose their effectiveness over time due to changes in market conditions and investor behavior 8279188 . This phenomenon is called Alpha Decay in quantitative finance. As a result, traders and investors must continuously search for new signals or develop new features that remain closely related to stock price movements. Staying adaptive and proactive is essential to maintain an edge in the dynamic and highly competitive financial markets.
@@
To deal with rapidly changing and complex financial markets, advanced models and techniques have been developed to manage the increasing volume and diversity of data. One of the most promising advances in recent years is the use of machine learning, particularly deep learning models, in the development of trading strategies MINTARYA202396 . These models can process large datasets and uncover complex patterns in market behavior that traditional methods might miss. However, despite their impressive capabilities, machine learning models still face challenges in areas such as interpretability Sejnowski_2020 , flexibility, and the need for extensive data labeling 10114634 ; kou2024automatestrategyfindingllm .
@@
A new development in this field involves the application of large language models (LLMs), such as OpenAI’s GPT models, for financial analysis. LLMs, which have been trained on large amounts of text data, can process financial news, reports, and historical price information to generate insights and predictions 10825946 . These models can generate a variety of outputs, including summarizing financial news and creating algorithmic strategies for trading. In particular, prompt-based LLMs, which take in specific prompts and generate relevant outputs based on those inputs, provide a powerful tool for automating and enhancing the alpha generation process in stock trading wang2024gptsignalgenerativeaisemiautomated . However, most of the time, other researchers only use LLMs to generate formulaic alphas and examine their correlations with stock return. They do not analyze how the LLM-generated alphas can be used in dynamic quant trading, which will be further investigated in this paper.
@@
In our previous study chen2025sentimentawarestockpriceprediction , LLM-generated formulaic alphas have been demonstrated to have strong predictive power in stock price prediction. In this study, we extend this line of research by shifting the focus from alpha generation to alpha integration and optimization. Instead of relying solely on the predictive strength of individual alphas, we investigate how reinforcement learning, specifically Proximal Policy Optimization, can be employed to dynamically allocate weights across multiple LLM-generated alphas. This approach enables the trading strategy to adapt to evolving market conditions, enhance robustness against noisy signals, and capture diverse market dynamics.
@@
First, we use a prompt-based LLM, deepseek-r1-distill-llama-70b model, to generate formulaic alphas for five major companies: Toyota, Apple, HSBC, Pepsi and Tencent. Alpha, in strict financial definition, refers to the excess return that a trading strategy can achieve above a benchmark index, such as the S&P 500 or Nikkei 225 kakushadze2016101formulaicalphas . In quantitative finance and machine learning, alpha factors often refer to signals or features that are predictive of future stock returns. Formulaic alpha involves mathematically defined trading signals or strategies aimed at predicting future asset returns. These strategies are usually expressed as explicit formulas that combine various types of data, including market, fundamental, or alternative inputs such as price, volume, volatility, and sentiment indicators. Traditionally, these formulas often originate from quantitative research. Equation 1 is an example of formulaic alpha. When a stock price is substantially higher than its recent average, equation 1 yields a negative value, indicating that the stock might be overbought and likely to decline. However, if the price is below the average, the alpha becomes positive, suggesting a potential buying opportunity due to mean reversion.
@@
PPO, a model-free algorithm proposed by schulman2017proximalpolicyoptimizationalgorithms , is particularly well-suited for continuous action spaces, making it ideal for determining the optimal weight distribution of different alphas in a portfolio. By training the PPO agent, this study seeks to identify the best combination of alphas that can maximize returns.
To evaluate the performance of our strategy, we compare the returns generated by the optimized alphas with those from an equal-weighted portfolio and major financial benchmarks, including the Nikkei 225, S&P 500, and Hang Seng Index. For example, the Nikkei 225 is one of the main stock market indices in Japan, tracking 225 large publicly traded companies in diverse sectors such as technology, automotive, finance and consumer goods. These benchmarks are widely used to represent the broader market and serve as reference points for assessing the effectiveness of trading strategies. In addition to cumulative returns, we also analyze the Sharpe ratio and maximum drawdown to provide a more comprehensive view of performance. The Sharpe ratio reflects the risk-adjusted return of the strategy, helping to assess whether higher returns are achieved efficiently. Maximum drawdown captures the largest peak-to-trough decline, indicating potential risk exposure during unfavorable market conditions. By evaluating these metrics together, we aim to determine not only whether LLM-generated alphas can outperform traditional market indices but also whether they can do so with more favorable risk-return characteristics.
@@
    Empirical analysis is conducted across multiple alpha selection strategies, including random, low-correlation, and high-contribution methods, demonstrating that the proposed framework consistently outperforms market benchmarks across different stocks. This provides practical insights into how alpha quality and selection criteria influence portfolio performance.
@@ 2 Related Work
Early efforts in formulaic alpha generation rely on statistical and econometric models. For example, Fama and French FAMA19933 propose factor models, which aim to explain asset returns using factors such as market risk, size, value, and momentum. These models have been foundational in understanding systematic risk factors, but they have limitations in capturing complex, non-linear relationships in financial data. More recent advancements have focused on enhancing traditional alpha models by incorporating machine learning algorithms.
@@
Machine learning approaches have gained significant attention in the domain of alpha generation. One notable development is the use of supervised learning models to predict stock prices and returns. Researchers have applied various machine learning techniques, including decision trees 8250694 , support vector machines (SVMs) 6703096 ; 9361804 , and random forests 9987903 , to identify patterns in historical data and generate alpha signals. However, these models are often limited by their reliance on structured data and the need for manual feature engineering, which can be time-consuming and prone to biases.
@@
A more recent shift in alpha generation research has involved the application of deep learning techniques, particularly convolutional neural networks (CNNs) 9317207 and recurrent neural networks (RNNs) 10392023 , to stock market prediction. These models are capable of automatically learning complex and non-linear patterns from raw market data. Recently, some researchers have proposed some hybrid models to further improve prediction performance. For example, Zhang et al. math11091985 propose a CNN-BiLSTM-Attention-based model to improve the performance of stock price forecasting. The proposed approach can outperform models such as LSTM. Lu et al. https://doi.org/10.1155/2020/6622927 propose a CNN-LSTM to predict the stock price.
@@
Recently, transformer-based models have become popular candidates in stock price forecasting. Some transformer models designed for time series forecasting include Dual Transformer chen2025sentiment , Informer zhou2021informerefficienttransformerlong , Autoformer wu2022autoformerdecompositiontransformersautocorrelation and TimeXer wang2024timexerempoweringtransformerstime . Li et al. li2023mastermarketguidedstocktransformer proposed a market-guided transformer to predict the stock price. This model uses momentary and cross-time stock correlation as well as market information for automatic feature selection.
@@
In addition to supervised learning and deep learning, natural language processing (NLP) techniques have been used to extract information from unstructured data sources such as financial news articles, earnings reports, and social media. NLP models, especially generative LLMs such as LLaMA, Mistral, and GPT, have shown promise in sentiment analysis and in predicting stock movements based on textual data 10825946 ; yang2023fingptopensourcefinanciallarge ; xie2023pixiulargelanguagemodel . These models analyze the sentiment of news and social media content and generate alpha signals based on the prevailing market sentiment. This approach has become more prominent due to the increasing availability of financial news and the growing importance of market sentiment in stock price movements.
@@
Prompt-based LLMs are another emerging area of interest in formulaic alpha generation. These models generate specific financial alphas based on customized input prompts. For example, researchers have explored the use of prompt-based LLMs to generate alphas that incorporate technical indicators, such as moving averages, RSI, and MACD, along with sentiment data from news articles or earnings calls kou2025automatestrategyfindingllm . For example, Chen and Kawashima chen2025sentimentawarestockpriceprediction utilize an LLM to generate formulaic alphas and provide evidence of their effectiveness in predicting stock price movements. Wang et al. wang2024gptsignalgenerativeaisemiautomated use the prompt-based GPT-4 model to generate formulaic alphas. By tailoring the prompts to include specific financial metrics, these LLMs can generate alphas that are directly aligned with the desired prediction task. This approach differs from traditional machine learning techniques because it combines domain-specific knowledge (in the form of technical indicators) with the generative capabilities of LLMs to create actionable trading strategies. However, their research only generates alphas but does not address how to adaptively combine them or assess their performance over time.
@@
Reinforcement learning (RL) has gained increasing attention in financial forecasting and trading. RL models are designed to learn optimal trading strategies by interacting with the market environment through trial and error. By continuously updating their policies based on observed rewards, RL algorithms can adapt to changing market dynamics and improve decision making over time Chen2025 . Various RL methods, such as Deep Q-Networks (DQN), Proximal Policy Optimization (PPO), and Soft Actor-Critic (SAC), have been successfully applied to tasks like portfolio management, asset allocation, and algorithmic trading huang2025deepreinforcementlearningframework ; ndikum2024advancinginvestmentfrontiersindustrygrade ; 10627674 ; ijfs11010010 . These approaches allow trading agents to make sequential decisions, balancing the trade-off between risk and return in complex and uncertain markets.
@@
The integration of LLMs and RL for optimizing multiple formulaic alphas is still a relatively novel area of research. Few studies have directly addressed the use of LLMs for alpha generation in combination with RL algorithms for weight optimization. We will examine the effectiveness of this framework in this study.
@@ 3.1 Data
Historical stock price data for Toyota, Apple, HSBC, Pepsi, and Tencent from 2016/02/16 to 2024/05/08 are downloaded using the yfinance Python library. This library provides an efficient way to access daily stock prices, including opening, closing, high, low, and volume data for each company. Stock data is used to calculate key technical indicators, such as moving averages, momentum, and RSI, which are incorporated into the alpha generation process. The Python library pandas-ta is used to calculate all the technical indicators. In finance, technical indicators, calculated from historical stock price and volume data, are often used by traders and investors to predict the stock trend and generate trading signals. Table 1 shows the technical indicators used in this study.
@@
The first 80% of the data will be used for training and the remaining 20% of the data will be used for testing.

### v2

@@ Abstract
This paper introduces a reinforcement learning framework that employs Proximal Policy Optimization (PPO) to dynamically optimize the weights of multiple large language model (LLM)-generated formulaic alphas for stock trading strategies. Formulaic alphas are mathematically defined trading signals derived from price, volume, sentiment, and other data. Although recent studies have shown that LLMs can generate diverse and effective alphas, a critical challenge lies in how to adaptively integrate them under varying market conditions. To address this gap, we leverage a DeepSeek model to generate fifty alphas for ten stocks, and then use PPO to adjust their weights in real time. Experimental results indicate that the PPO-optimized strategy does not consistently deliver the highest cumulative returns across all stocks, but it achieves comparatively higher Sharpe ratios and smaller maximum drawdowns in most cases. When compared with baseline strategies, including equal-weighted, buy-and-hold, random entry/exit, and momentum approaches, PPO demonstrates more stable risk-adjusted performance. The findings highlight the importance of reinforcement learning in the allocation of alpha weights and show the potential of combining LLM-generated signals with adaptive optimization for robust financial forecasting and trading.
@@ 1 Introduction
Over the past few decades, financial markets have evolved significantly with the integration of advanced technologies, especially in quantitative finance. Traditionally, stock trading strategies have relied on financial indicators such as moving averages, volatility measures, and momentum indicators to predict market behavior and generate profits Pardeshi and Kale (2021); Mostafavi and Hooman (2025). These methods have been successful to some extent, but often lack the flexibility and adaptability required to respond to changing market conditions. In addition, relying solely on manually chosen indicators may restrict the discovery of new patterns within the extensive and constantly changing financial data landscape.
@@
Furthermore, trading signals often lose their effectiveness over time due to changes in market conditions and investor behavior Zhou and Lin (2017). This phenomenon is called Alpha Decay in quantitative finance. As a result, traders and investors must continuously search for new signals or develop new features that remain closely related to stock price movements. Staying adaptive and proactive is essential to maintain an edge in the dynamic and highly competitive financial markets.
@@
To deal with rapidly changing and complex financial markets, advanced models and techniques have been developed to manage the increasing volume and diversity of data. One of the most promising advances in recent years is the use of machine learning, particularly deep learning models, in the development of trading strategies Mintarya et al. (2023). These models can process large datasets and uncover complex patterns in market behavior that traditional methods might miss. However, despite their impressive capabilities, machine learning models still face challenges in areas such as interpretability Sejnowski (2020), flexibility, and the need for extensive data labeling Gao and Guan (2023); Kou et al. (2024).
@@
A new development in this field involves the application of large language models (LLMs), such as OpenAI’s GPT models, for financial analysis Chen (2025c). LLMs, which have been trained on large amounts of text data, can process financial news, reports, and historical price information to generate insights and predictions Chen and Kawashima (2024). These models can generate a variety of outputs, including summarizing financial news and creating algorithmic strategies for trading. In particular, prompt-based LLMs, which take in specific prompts and generate relevant outputs based on those inputs, provide a powerful tool for automating and enhancing the alpha generation process in stock trading Wang et al. (2024a). However, most of the time, other researchers only use LLMs to generate formulaic alphas and examine their correlations with stock return. They do not analyze how the LLM-generated alphas can be used in dynamic quant trading, which will be further investigated in this paper.
@@
In our previous study Chen and Kawashima (2025a), LLM-generated formulaic alphas have been demonstrated to have strong predictive power in stock price prediction. In this study, we extend this line of research by shifting the focus from alpha generation to alpha integration and optimization. Instead of relying solely on the predictive strength of individual alphas, we investigate how reinforcement learning, specifically Proximal Policy Optimization, can be employed to dynamically allocate weights across multiple LLM-generated alphas. This approach enables the trading strategy to adapt to evolving market conditions, enhance robustness against noisy signals, and capture diverse market dynamics.
@@
First, we use a prompt-based LLM, deepseek-r1-distill-llama-70b model, to generate formulaic alphas for ten major companies: Toyota, Apple, HSBC, Pepsi, Tencent, Airbus, Exxon Mobil, Petrobras, Netfilx, and InfuSystem. Alpha, in strict financial definition, refers to the excess return that a trading strategy can achieve above a benchmark index, such as the S&P 500 or Nikkei 225 Kakushadze (2016). In quantitative finance and machine learning, alpha factors often refer to signals or features that are predictive of future stock returns. Formulaic alpha involves mathematically defined trading signals or strategies aimed at predicting future asset returns. These strategies are usually expressed as explicit formulas that combine various types of data, including market, fundamental, or alternative inputs such as price, volume, volatility, and sentiment indicators. Traditionally, these formulas often originate from quantitative research. Equation 1 is an example of formulaic alpha. When a stock price is substantially higher than its recent average, Equation 1 yields a negative value, indicating that the stock might be overbought and likely to decline. However, if the price is below the average, the alpha becomes positive, suggesting a potential buying opportunity due to mean reversion.
@@
PPO, a model-free algorithm proposed by Schulman et al. (2017), is particularly well-suited for continuous action spaces, making it ideal for determining the optimal weight distribution of different alphas in a portfolio. Unlike off-policy methods such as Deep Deterministic Policy Gradient (DDPG) and Soft Actor-Critic (SAC), which are often sensitive to hyperparameter choices and can struggle in non-stationary environments, PPO employs a clipped surrogate objective function to regulate policy updates. This design effectively limits the magnitude of each update, acting as a trust-region–like constraint that helps avoid abrupt changes in the trading strategy, which are particularly problematic under the low signal-to-noise conditions of financial markets Lin and Beling (2021). Moreover, by adopting an on-policy learning framework, PPO promotes more stable learning dynamics Zhang et al. (2022) and a balanced exploration–exploitation behavior, reducing the tendency to adapt excessively to short-term market noise. For example, Rio et al. Rio et al. (2024) show that PPO has a more stable training process compared to Asynchronous Advantage Actor-Critic (A3C). Mammadzada Mammadzada (2025) finds that PPO has more stable and consistent performance compared to SAC. Consequently, by training the PPO agent, this study seeks to identify a robust combination of alphas that maximizes risk-adjusted returns while ensuring training stability across varying market regimes.
To evaluate the performance of our strategy, we compare its returns with those of several baseline approaches, including an equal-weighted portfolio, a buy-and-hold strategy, a random entry/exit strategy, and a momentum strategy. In addition to cumulative returns, we also analyze the Sharpe ratio and maximum drawdown to provide a more comprehensive view of performance. The Sharpe ratio reflects the risk-adjusted return of the strategy, helping to assess whether higher returns are achieved efficiently. Maximum drawdown captures the largest peak-to-trough decline, indicating potential risk exposure during unfavorable market conditions. By evaluating these metrics together, we aim to determine not only whether LLM-generated alphas can outperform traditional market indices but also whether they can do so with more favorable risk-return characteristics.
@@
    Through an ablation study comparing human-crafted and LLM-generated alphas, this study demonstrates the superior performance of LLM-generated alphas.
  3. 3.
    Empirical analysis is conducted across multiple alpha selection strategies, including random, low-correlation, and high-contribution methods, demonstrating the consistent performance of the proposed framework across different stocks. This provides practical insights into how alpha quality and selection criteria influence portfolio performance.
@@ 2 Related Work
Early efforts in formulaic alpha generation rely on statistical and econometric models. For example, Fama and French Fama and French (1993) propose factor models, which aim to explain asset returns using factors such as market risk, size, value, and momentum. These models have been foundational in understanding systematic risk factors, but they have limitations in capturing complex, non-linear relationships in financial data. More recent advancements have focused on enhancing traditional alpha models by incorporating machine learning algorithms.
@@
Machine learning approaches have gained significant attention in the domain of alpha generation. One notable development is the use of supervised learning models to predict stock prices and returns. Researchers have applied various machine learning techniques, including decision trees Kamble (2017), support vector machines (SVMs) Hu et al. (2013); Liu et al. (2020), and random forests Du et al. (2022), to identify patterns in historical data and generate alpha signals. However, these models are often limited by their reliance on structured data and the need for manual feature engineering, which can be time-consuming and prone to biases Chen (2025f).
@@
A more recent shift in alpha generation research has involved the application of deep learning techniques, particularly convolutional neural networks (CNNs) Mehtab and Sen (2020); Chen (2025d) and recurrent neural networks (RNNs) Shinde et al. (2023), to stock market prediction. These models are capable of automatically learning complex and non-linear patterns from raw market data. Recently, some researchers have proposed some hybrid models to further improve prediction performance. For example, Zhang et al. Zhang et al. (2023) propose a CNN-BiLSTM-Attention-based model to improve the performance of stock price forecasting. The proposed approach can outperform models such as LSTM. Lu et al. Lu et al. (2020) propose a CNN-LSTM to predict the stock price.
@@
Recently, transformer-based models have become popular candidates in stock price forecasting Chen (2025b). Some transformer models designed for time series forecasting include Dual Transformer Chen and Kawashima (2025b), Informer Zhou et al. (2021), Autoformer Wu et al. (2022) and TimeXer Wang et al. (2024b). For example, Li et al. Li et al. (2023) proposed a market-guided transformer to predict the stock price. This model uses momentary and cross-time stock correlation as well as market information for automatic feature selection.
@@
In addition to supervised learning and deep learning, natural language processing (NLP) techniques have been used to extract information from unstructured data sources such as financial news articles, earnings reports, and social media Chen (2025a). NLP models, especially generative LLMs such as LLaMA, Mistral, and GPT, have shown promise in sentiment analysis and in predicting stock movements based on textual data Chen and Kawashima (2024); Yang et al. (2023); Xie et al. (2023). These models analyze the sentiment of news and social media content and generate alpha signals based on the prevailing market sentiment. This approach has become more prominent due to the increasing availability of financial news and the growing importance of market sentiment in stock price movements.
@@
Prompt-based LLMs are another emerging area of interest in formulaic alpha generation. These models generate specific financial alphas based on customized input prompts. For example, researchers have explored the use of prompt-based LLMs to generate alphas that incorporate technical indicators, such as moving averages, RSI, and MACD, along with sentiment data from news articles or earnings calls Kou et al. (2025). For example, Chen and Kawashima Chen and Kawashima (2025a) utilize an LLM to generate formulaic alphas and provide evidence of their effectiveness in predicting stock price movements. Wang et al. Wang et al. (2024a) use the prompt-based GPT-4 model to generate formulaic alphas. By tailoring the prompts to include specific financial metrics, these LLMs can generate alphas that are directly aligned with the desired prediction task. This approach differs from traditional machine learning techniques because it combines domain-specific knowledge (in the form of technical indicators) with the generative capabilities of LLMs to create actionable trading strategies. However, their research only generates alphas but does not address how to adaptively combine them or assess their performance over time.
@@
Reinforcement learning (RL) has gained increasing attention in financial forecasting and trading. RL models are designed to learn optimal trading strategies by interacting with the market environment through trial and error. By continuously updating their policies based on observed rewards, RL algorithms can adapt to changing market dynamics and improve decision making over time Chen (2025e). Various RL methods, such as Deep Q-Networks (DQN), Proximal Policy Optimization (PPO), and Soft Actor-Critic (SAC), have been successfully applied to tasks like portfolio management, asset allocation, and algorithmic trading Huang et al. (2025); Ndikum and Ndikum (2024); Sharma et al. (2024); Millea and Edalat (2023). These approaches allow trading agents to make sequential decisions, balancing the trade-off between risk and return in complex and uncertain markets. For example, Lee and Moon Lee and Moon (2023) propose a transformer-based actor–critic reinforcement learning model that uses past stock price information to guide trading decisions and shows improved risk-adjusted returns across multiple stock market datasets compared with existing methods. Orra et al. Orra et al. (2025) introduce an expert-guided reward design for deep reinforcement learning in stock trading, where technical indicator signals provide feedback to stabilize learning and improve trading performance across multiple global stock markets. Kong and So Kong and So (2023) extend ensemble-based reinforcement learning for stock trading by combining multiple actor–critic methods into a larger ensemble and applying it across several global stock markets, showing more stable returns in empirical tests. Guevara Guevara (2025) proposes an actor–critic reinforcement learning trading model with an adaptive data window, showing reduced losses and faster learning across crude oil, gold, and Euro markets. Sattar et al. Sattar et al. (2025) propose a reinforcement learning trading framework that combines price data, sentiment indicators, and drawdown-based rewards, showing improved risk control and returns compared with baseline market strategies.
@@
Although reinforcement learning has been extensively applied in finance, the integration of LLMs and RL for optimizing multiple formulaic alphas is still a relatively novel area of research. Few studies have directly addressed the use of LLMs for alpha generation in combination with RL algorithms for weight optimization. We will examine the effectiveness of this framework in this study.
@@ 3.1 Data
Since multiple related studies Vijh et al. (2020); Yuan et al. (2020); Fu et al. (2023); GE et al. (2022) in stock forecasting or trading typically focus on one to ten stocks, this study selects ten stocks to balance representativeness and computational cost. The selected stocks—Toyota, Apple, HSBC, Pepsi, Tencent, Airbus, Exxon Mobil, Netflix, Petrobras, and InfuSystem—come from different regions, including the U.S., Europe, China, Japan, and Brazil, and cover a broad range of industries such as automotive, technology, finance, energy, aerospace, healthcare services, and consumer goods. This diversity helps evaluate whether the proposed method works across different market conditions rather than for a single sector. Historical stock price data from 2016/02/16 to 2024/05/08 are obtained using the yfinance Python library. This library provides an efficient way to access daily stock prices, including opening, closing, high, low, and volume data for each company.
In addition, stock data is used to calculate key technical indicators, such as moving averages, momentum, and RSI, which are incorporated into the alpha generation process. The Python library pandas-ta is used to calculate all the technical indicators. In finance, technical indicators, calculated from historical stock price and volume data, are often used by traders and investors to predict the stock trend and generate trading signals. Table 1 shows the technical indicators used in this study.
@@
The first 80% of the data will be used for training and the remaining 20% of the data will be used for testing. We adopt an 80/20 split between the training and test sets to achieve a balanced trade-off between model learning and out-of-sample evaluation. Prior work by Joseph (2022) suggests an optimal test proportion of approximately $1/(\sqrt{p}+1)$. In our setting, where $p=50$ alpha features are used as predictors, this guideline corresponds to a test size of about 12.3%.
We deliberately choose a larger test split of 20% for a practical reason. From a statistical standpoint, a larger hold-out period enables a more reliable evaluation of performance metrics such as the Sharpe Ratio and Maximum Drawdown, reducing the likelihood that the results are driven by short-term market fluctuations.
In Section 5.3, we also conduct a walk-forward optimization analysis to further examine the robustness of the reinforcement learning framework.